OpenStudy (anonymous):

…square root of 15 Times the square root of 48 write your answer in radical notation

OpenStudy (anonymous):

Divide it into parts : Square root of 15 that is \(\sqrt{15}\) Times(Multiply) the square root of 48 that is : \(\times \sqrt{48}\) So, finally you will get : \[\sqrt{15} \times \sqrt{48} = Solve \; \; this..\]

OpenStudy (anonymous):

Tell me : \[\sqrt{15} = ??\]

OpenStudy (anonymous):

I need to know what those together equal. I know it's square root of 720but I need the radical notation

OpenStudy (anonymous):

 start from here